150 Literaturverzeichnis

[57] V. I. Arnol’d, Mathematical Methods of Classical Mechanics, Band 60 der Reihe Graduate Texts in Mathematics (Springer, New York, 1984).

[58] M. V. Berry und C. Upstill, in Progress in Optics, herausgegeben von E. Wolf (North-Holland, Amsterdam, 1980), Band VIII, S. 257–346.

[59] D. Ruelle, Ann. Inst. H. Poincaré 42, 109 (1985).

[60] U. Dressler und W. Lautenborn, Phys. Rev. A 41, 6702 (1990).

[61] B. Eckhardt und D. Wintgen, J. Phys. A 24, 4335 (1991).

[62] J. M. Robbins und R. G. Littlejohn, Phys. Rev. Lett. 58, 1388 (1987).

[63] J. M. Robbins, Nonlinearity 4, 343 (1991).

[64] J. Robbins, Chaos 2, 145 (1992).

[65] J. P. Eckmann und R. Sénéor, Arch. Rat. Mech. Anal. 61, 153 (1976).

[66] I. C. Percival, Adv. Chem. Phys. 36, 1 (1977).

[67] V. P. Maslov und M. V. Fedoryuk, Semiclassical Approximations in Quantum Mechanics (Reidel, Dordrecht, 1981).

[68] S. C. Creagh, J. M. Robbins und R. G. Littlejohn, Phys. Rev. A 42, 1907 (1990).

[69] M. V. Berry, Proc. R. Soc. Lond. A 392, 45 (1984).

[70] J. H. Hannay, J. Phys. A 18, 221 (1985).

[71] R. J. Rimmer, Memoirs of the AMS 272 (American Mathematical Society, Providence, Rhode Island, 1983).

[72] M. Golubitsky und I. Stewart, Physica D 24, 391 (1987).

[73] M. A. M. de Aguiar, C. P. Malta, M. Baranger und K. T. R. Davies, Ann. Phys. (N. Y.) 180, 167 (1987).

[74] M. A. M. de Aguiar und C. P. Malta, Physica D 30, 413 (1988).

[75] A. M. Ozorio de Almeida und M. A. M. de Aguiar, Physica D 41, 391 (1990).

[76] G. D. Birkhoff, Dynamical Systems (American Mathematical Society, New York, 1927).

[77] L. E. Reichl, The Transition to Chaos (Springer, New York, 1992).

[78] J. Main und G. Wunner, Phys. Rev. A 55, 1743 (1997).

[79] A. Mostowski und M. Stark, Introduction to Higher Algebra (Pergamon Press, Oxford, 1964).

[80] M. V. Berry und N. L. Balasz, J. Phys. A 12, 625 (1979).

[81] E. J. Heller, S. Tomsovic und M. A. Sepúlveda, Chaos 2, 105 (1992).

[82] I. S. Gradshteyn und I. M. Rhyzik, Table of Integrals, Series, and Products (Academic Press, San Diego, 1994).

[83] F. J. Wright, J. Phys. A 13, 2913 (1980).

[84] T. Pearcey, Philos. Mag. 37, 311 (1946).

[85] S. Tomsovic, M. Grinberg und D. Ullmo, Phys. Rev. Lett. 75, 4346 (1995).

[86] D. Ullmo, M. Grinberg und S. Tomsovic, Phys. Rev. E 54, 136 (1996).

[87] D. Wintgen, Phys. Rev. Lett. 58, 1589 (1987).

[88] U. Eichmann, K. Richter, D. Wintgen und W. Sandner, Phys. Rev. Lett. 61, 2438 (1988).